OpenStudy (anonymous):

a studen opens a book to two facing pages. the product of the page numbers is 1332 find the page numbers? a studen opens a book to two facing pages. the product of the page numbers is 1332 find the page numbers? @Mathematics

OpenStudy (across):

This question can be reworded to something like this: "What are two consecutive numbers whose product is 1332?"

OpenStudy (across):

\[x(x+1)=1332\]\[x^2+x=1332\]\[x^2+x+\frac{1}{4}=1332+\frac{1}{4}\]\[\left ( x+\frac{1}{2} \right )^2=\frac{5329}{4}\]\[x+\frac{1}{2}=\pm\frac{73}{2}\]\[x=36,-37\]

OpenStudy (across):

You pick the page numbers. ;)

OpenStudy (asnaseer):

@Zarkon has a neat way of solving these types of problems, he just finds the factors of the product in question and "sees" the solution emerge - works on most of these types of problems. e.g.: \[1332=2^2*3^2*37=4*9*37=36*37\]so the page numbers are 36 and 37.